OpenStudy (brooke2353):

Will Give medal!! In a certain country ½ of 5 = 3. If the same proportion holds, what is the value of 1/3 of 10 ?

OpenStudy (anonymous):

4

OpenStudy (brooke2353):

explain your answer please

OpenStudy (anonymous):

5/2=3 and 10/3 = x (10/3)/(5/2) = 20/15 = 4/3 So x/3 = 4/3 x=4

OpenStudy (brooke2353):

thanks